### v3

@@ IV Future opportunities
Challenge scenarios and real-world studies Navigating crowded environments is a significant challenge for the visually impaired. However, studies included in the SLR that address this issue are limited. Among the studies that considered this challenge, evaluations were conducted in controlled settings, rather than addressing the complexities of real-world scenarios. It is essential to note that the applicability of these solutions to real-world scenarios such as train stations is crucial. Therefore, there is a pressing need for future studies focusing on the complexities of navigation in high-traffic public places. Similarly, confronting challenging conditions, such as changes in illumination, low-light scenarios, high-speed dynamic objects, and complex backgrounds, can enhance the robustness and versatility of navigation systems. To overcome these challenges in the field of visually impaired navigation, we can draw inspiration from other domains and adapt the proposed solutions. For instance, [116] introduced a system that utilized an image enhancement technique for ORB points and LSD line feature recovery. This approach enables the system to perform effectively in a broader range of scenarios, providing extensive spatial information, even in low-light and hazy agricultural environments.
@@
Long-term navigation The development of solutions that are effective over extended navigation periods is critical for achieving autonomous navigation. These solutions must ensure accurate mapping and localization even when the maps are updated over a longer navigation duration. To address this challenge, researchers can leverage the proposed solutions in robotics, such as that presented in [117], which introduced a novel long-term SLAM system with map prediction and dynamic removal, allowing wheelchair robots to maintain precise navigation capabilities over extended periods.
@@
Deep learning integration The integration of deep learning with the SLAM algorithms, adopted for BVI navigation, deserves further attention. Deep learning offers a versatile approach for enhancing various aspects of SLAM, such as precise pose estimation under challenging conditions, relocalization, and loop-closure detection. Despite challenges, such as the need for large, accurately labeled datasets, the black-box nature of deep-learning models, and the computational intensity of SLAM systems, the association between deep learning and SLAM holds promise for advancing navigation solutions for the visually impaired, particularly in challenging scenarios.
Indoor and outdoor navigation integration The need for navigation systems capable of seamless transitions in both indoor and outdoor environments is crucial for enhancing the independence and mobility of BVI individuals. However, most of the reviewed papers in our SLR have primarily concentrated on indoor environments. Future research could address this challenge by developing solutions that offer unified and consistent navigation experience.
Obstacle detection Achieving detailed knowledge of obstacles and their characteristics is essential for BVI people. Although some papers included in the SLR addressed obstacle detection, the depth and accuracy of the obstacle information provided may still be limited. Future research opportunities lie in enhancing the SLAM algorithms to provide context-aware obstacle information. Researchers can draw inspiration from approaches adopted in robotics and autonomous drones such as real-time metric-semantic SLAM, as demonstrated in [118].
@@
Semantic information integration Integrating semantic information into SLAM algorithms can significantly enhance the performance and robustness of navigation systems for BVI individuals. This information can be used to refine the mapping and localization processes and enhance the overall reliability of navigation in complex environments. For instance, semantic information can help reject outliers in loop closure detection, which is a crucial step in SLAM that involves identifying and matching previously visited locations.
@@
Product development and collaboration Notably, all the reviewed approaches were prototypes in the early stages of research and are not yet practical products. This might be due to the absence of a unified community or group dedicated to solving the BVI navigation challenges. Much of the work in this domain is carried out by academic groups or small companies that often fail to produce feasible final products. This underscores a significant future opportunity to develop collaboration and to bridge the gap between research and practical implementation.
@@ V Conclusion
This study presents a systematic literature review of recent studies on SLAM-based solutions for BVI navigation. Excluding papers published before 2017, this review focused on the latest advancements, innovations, and considerations, resulting in a more relevant and comprehensive understanding of the current state of research. Relevant data were extracted from 54 selected studies that adhered to the SLR selection criteria to address the research questions. By analyzing the selected papers based on their SLAM techniques, we observed that the majority of the studies utilized visual SLAM techniques, such as ORB-SLAM3, owing to their advantages for visual sensors.
@@
We also investigated the challenging scenarios encountered by SLAM-based navigation systems, which have been addressed in the literature. Furthermore, we discuss how the proposed SLAM-based solutions are expected to improve the mobility and navigation of the visually impaired. Finally, we identified gaps, opportunities, and areas of interest that could be explored further in future research, such as addressing challenges in crowded environments, improving real-world applicability, integrating deep learning, and ensuring long-term navigation effectiveness in SLAM-based solutions for visually impaired navigations.

### v4

@@ IV Future opportunities
Challenge scenarios and real-world studies Navigating crowded environments remains a significant challenge for the visually impaired and studies addressing this issue are limited. Evaluations are often conducted in controlled settings rather than real-world scenarios. Future research should focus on developing and testing solutions in high-traffic public places such as train stations and shopping malls. Furthermore, addressing challenging conditions, such as changes in illumination, low-light scenarios, high-speed dynamic objects, and complex backgrounds, can enhance the robustness and versatility of navigation systems. Techniques such as image enhancement for ORB points and LSD line feature recovery used in agricultural environments [116] can be adapted for visually impaired navigation.
Future research should prioritize real-world testing and validation by conducting extensive field tests in high-traffic public areas, such as train stations, airports, and shopping centers, to evaluate the system performance under diverse environmental factors. Handling dynamic and complex environments requires the development of algorithms that can adapt to rapid changes such as sudden variations in lighting, moving objects, and different background textures. Machine learning models that can learn and predict common patterns in dynamic environments can improve system responsiveness and reliability.
To enhance robustness under low-light and variable lighting conditions, one solution is to integrate advanced image enhancement techniques and utilize sensors complementary to traditional cameras, ensuring better obstacle detection and navigation assistance in poorly lit areas. It is essential that these enhancements do not compromise the real-time performance of the system and maintain seamless and immediate feedback for the user.
@@
Long-term navigation The development of solutions that are effective over extended navigation periods is critical to achieve autonomous navigation. These solutions must ensure accurate mapping and localization even when the maps are updated over a longer navigation duration. To address this challenge, researchers can leverage the proposed solutions in robotics, such as that presented in [117], which introduced a novel long-term SLAM system with map prediction and dynamic removal, thereby allowing wheelchair robots to maintain precise navigation capabilities over extended periods. Future research should focus on developing robust algorithms for continuous map updates and maintenance, including strategies for handling environmental changes over time. This involves dynamic object removal and adaptive map refinement to ensure the accuracy and relevance of navigation maps.
@@
Deep learning integration The integration of deep learning with SLAM algorithms for BVI navigation requires further investigation. Deep learning offers a versatile approach for enhancing various aspects of SLAM such as precise pose estimation under challenging conditions, relocalization, and loop-closure detection. Despite challenges, such as the need for large, accurately labeled datasets, the black-box nature of deep-learning models, and the computational intensity of SLAM systems, the association between deep learning and SLAM holds promise for advancing navigation solutions for the visually impaired, particularly in challenging scenarios. Future research should focus on developing more efficient deep-learning models that can operate effectively with limited computational resources and real-time constraints. In addition, creating large-scale, accurately labeled datasets tailored for BVI navigation is crucial for training robust models. Addressing the interpretability of deep learning models can also enhance the trust and transparency in these systems. Collaboration among machine-learning experts, roboticists, and vision scientists can drive the development of innovative algorithms that leverage deep learning to enhance the reliability and accuracy of SLAM-based navigation aids for the visually impaired.
Indoor and outdoor navigation integration Seamless transitions between indoor and outdoor environments are crucial for enhancing the independence and mobility of BVI individuals. However, most studies in our SLR have focused primarily on indoor environments. Future research should aim to develop solutions that provide unified and consistent navigation experience in both indoor and outdoor settings. Researchers should explore the integration of robust sensor fusion techniques and adaptive algorithms capable of handling the different conditions and challenges of these environments. Additionally, extensive real-world testing in various settings is essential to ensure the practicality and reliability of these solutions.
Obstacle detection Achieving detailed knowledge of obstacles and their characteristics is essential for BVI people. Although some studies included in the SLR addressed obstacle detection, the depth and accuracy of the obstacle information provided may still be limited. To address the need for more detailed and accurate obstacle information for BVI individuals, future research should focus on advancing SLAM algorithms to deliver context-aware obstacle detection. This involves integrating semantic understanding with precise spatial measurements, allowing the system to identify and interpret the nature and significance of obstacles accurately.
Researchers should explore methods to enhance the real-time processing capabilities to ensure that the system can operate effectively in dynamic and complex environments. Additionally, it is crucial to develop algorithms that can learn and adapt to various obstacle types and scenarios through continuous data collection and machine learning.
Drawing inspiration from approaches used in robotics and autonomous drones, such as the real-time metric-semantic SLAM demonstrated by [118], can provide valuable insights. This approach integrates semantic understanding with precise distance and spatial relationship measurements, enabling more comprehensive and reliable navigation aids. Therefore, future research should prioritize improving both the depth and accuracy of obstacle information, while ensuring robust real-time performance and adaptability to various real-world conditions.
@@
Semantic information integration Integrating semantic information into SLAM algorithms can significantly enhance the performance and robustness of navigation systems for BVI individuals. This information can be used to refine the mapping and localization processes and enhance the overall reliability of navigation in complex environments. For instance, semantic information can help reject outliers in loop closure detection, which is a crucial step in SLAM, which involves identifying and matching previously visited locations.
To advance this area, future research should focus on developing advanced techniques for semantic data extraction and integration within the SLAM frameworks. Researchers should also explore methods to ensure real-time performance while maintaining the accuracy and detail of the semantic information.
@@
Product development and collaboration Notably, all the reviewed approaches were prototypes in the early stages of research and are not yet practical. This might be due to the absence of a unified community or group dedicated to solving the BVI navigation challenges. Much of the work in this domain has been carried out by academic groups or small companies that often fail to produce feasible final products. This underscores a significant future opportunity to develop collaboration and to bridge the gap between research and practical implementation. Additionally, efforts should be made to develop standardized evaluation metrics and protocols to ensure that the developed systems meet real-world needs and can be effectively transitioned from prototypes to market-ready solutions. Encouraging partnerships with technology companies can also accelerate the commercialization process. These partnerships provide the necessary support to bring innovative solutions to the market.
@@ V Conclusion
This study presents a systematic literature review of recent studies on SLAM-based solutions for BVI navigation. Excluding papers published before 2017, this review focused on the latest advancements, innovations, and considerations, resulting in a more relevant and comprehensive understanding of the current state of research. The insights provided by this systematic literature review are intended to guide researchers in the academic and research communities. They inform the existing gaps and future opportunities to address the challenges faced by SLAM-based assistive solutions.
Relevant data were extracted from 54 selected studies that adhered to the SLR selection criteria to address the research questions. By analyzing the selected papers based on their SLAM techniques, we observed that the majority of the studies utilized visual SLAM techniques, such as ORB-SLAM3, owing to their advantages for visual sensors.
@@
We investigated the challenging scenarios encountered by SLAM-based navigation systems, which have been addressed in the literature. Additionally, we discussed practical considerations that affect the usability and adoption of these systems. Furthermore, we analyzed how the proposed SLAM-based solutions improve the mobility and navigation of visually impaired individuals. We evaluated the effectiveness of these solutions in real-world scenarios and assessed the user satisfaction to understand their practical impact on BVI mobility. Finally, we identified gaps, opportunities, and areas of interest that could be explored further in future research, such as addressing challenges in crowded environments, improving real-world applicability, integrating deep learning, and ensuring long-term navigation effectiveness in SLAM-based solutions for visually impaired navigations.